OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL FOR ANSWER 1. Solve for y: (2/5)y -7= 23

OpenStudy (danjs):

2y -35 = 23*5

OpenStudy (danjs):

y = (23*5 +35) / 2

OpenStudy (anonymous):

Can u show me steps how u got 35 like i dont understand inequalities can u plz walk me through or show me your work on how u did the problem plz

OpenStudy (danjs):

multiply everything by 5 to get rid of the fraction

OpenStudy (danjs):

\[5*[\frac{ 2 }{ 5 }y - 7] = 5*[23]\]

OpenStudy (danjs):

2y - (5*7) = 5*23

OpenStudy (danjs):

2y - 35 = 115

OpenStudy (danjs):

2y = 150 y = 75

OpenStudy (anonymous):

how did u get 2y=150 y=75

OpenStudy (anonymous):

please help me @DanJS

OpenStudy (danjs):

hi

OpenStudy (anonymous):

how did you get 2y=150 y=75

OpenStudy (danjs):

2 y - 35 = 115 add 35 to both sides of the equation

OpenStudy (danjs):

2y -35 + 35 = 115 + 35

OpenStudy (danjs):

2y = 150

OpenStudy (anonymous):

how u get y=75

OpenStudy (danjs):

divide both sides by 2 \[\frac{ 2 }{ 2 }y = \frac{ 150 }{ 2 }\]

OpenStudy (danjs):

1y = 75

OpenStudy (anonymous):

2. Determine the type of boundary line and shading for 2x + 4y greater than or equal to 10 @DanJS

OpenStudy (danjs):

do you have to graph it?

OpenStudy (anonymous):

it doesnt say :(((

OpenStudy (anonymous):

i think determine if it is dashed or something i have NO idea :(

OpenStudy (danjs):

ok, if it is greater than or equal to, it is a solid line If it is just greater than or less than it is a dashed line

OpenStudy (danjs):

Dashed Line \[\gt or \lt \]

OpenStudy (danjs):

Solid Line \[\le or \ge\]

OpenStudy (anonymous):

2x+4y_<10 thats how i meant to write the 1 part not greater to or equal to i think

OpenStudy (danjs):

if y is greater than or equal, you shade above the line If y is Less than or equal to, you shade below the line

OpenStudy (danjs):

ok \[4y \lt -2x + 10\] \[y \lt \frac{ -2 }{ 4 }x + \frac{ 10 }{ 4 }\]

OpenStudy (danjs):

The line is dashed and is shaded below

OpenStudy (danjs):